4.60 × 10⁻³ kg of sodium chloride was dissolved in 0.130 kg of water at 280 K. The osmotic pressure of the solution was found to be 2.66 × 10⁶ Pa. To calculate the van't Hoff factor and the degree of dissociation. The density of the solution is 1.02 × 10³ kg m⁻³ and the molar mass of NaCl is 58.5 × 10⁻³ kg mol⁻¹.
1
Expert's answer
2020-12-25T05:11:27-0500

Osmotic pressure = i . C . R . T


2.66 × 106 =

i × ( 4.60 × 1000 / 130) × 8.314 × 280.


i = 2.66 ×106 ×130 / 4.60 × 1000 × 8.314 × 280


i = 2 ( approx ) .


So, 100 % ionise ( approx).